# Towards a Benchmark for Causal Business Process Reasoning with LLMs

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used for boosting organizational efficiency and automating tasks. While not originally designed for complex cognitive processes, recent efforts have further extended to employ LLMs in activities such as reasoning, planning, and decision-making. In business processes, such abilities could be invaluable for leveraging on the massive corpora LLMs have been trained on for gaining deep understanding of such processes. In this work, we plant the seeds for the development of a benchmark to assess the ability of LLMs to reason about causal and process perspectives of business operations. We refer to this view as Causally-augmented Business Processes (BPC). The core of the benchmark comprises a set of BPC related situations, a set of questions about these situations, and a set of deductive rules employed to systematically resolve the ground truth answers to these questions. Also with the power of LLMs, the seed is then instantiated into a larger-scale set of domain-specific situations and questions. Reasoning on BPC is of crucial importance for process interventions and process improvement. Our benchmark could be used in one of two possible modalities: testing the performance of any target LLM and training an LLM to advance its capability to reason about BPC.
@@
While reasoning may not be inherent in the architecture of LLMs, critics do argue that the ability to reason does manifest itself as an emergent property in LLMs [20]. That is, once trained on sufficiently large corpora of examples, LLMs may become capable of statistically deriving statements that are also coincidentally sound with the given set of arguments in the input. Similar emergent abilities may also include planning, decision-making, in-context learning, and answering in zero-shot settings [15]. Whether reasoning is an inherent capability or an emergent property of future LLMs, particularly for interpreting business process models, is likely to remain a heated debate. This debate emphasizes the importance of establishing benchmarks for evaluating LLM performance. Our focus is on driving such enablement by using LLMs to analyze textual descriptions of Causally-augmented Business Processes (BPCs). This refers to a specific type of business process where causal execution dependencies among the activities are a first-class citizen of the process descriptions. Teaching LLMs to reason about BPCs could enhance the analysis and improvement of such processes.
To standardize and measure the ability of LLMs to facilitate such tasks, a designated benchmark is being developed. The purpose of the developed novel benchmark is twofold: as a testing dataset that can be employed to quantify the ability of an LLM to reason about BPCs, and as a training dataset that can be used to adapt the ability of an LLM for this task. Our developed prototype is intended to gradually evolve into a full-scale benchmark. We evaluated two open-source and three commercial LLMs on a small subset of the situations and questions. Our results highlight the importance of producing an objective numeric scale partitioned by different perspectives to compare and assess various LLM performances. They also indicate that there is room for further improvement of LLMs in the task of reasoning about BPCs. The evolving benchmark and corresponding materials are available here: https://github.com/IBM/SAX/tree/main/NLP4BPM2024.
@@ 2 Background
The potential of leveraging LLMs in the BPM field has been recently researched, for example, by analyzing which opportunities and challenges LLMs pose for the individual stages of the BPM lifecycle [19], in [4], where prompt engineering techniques are discussed as an alternative to fine-tuning a specific LLM, and in [12] where the authors introduce the notion of a Large Process Model (LPM), an envisioned neuro-symbolic software system that integrates process management knowledge in organizations with LLMs and statistical and inference methods for the automated inference of insights and actions. In addition, LLMs have been researched in a wide range of tasks related to BPM, including process mining [9, 3], automation of portions of complex tasks 222https://www.infoworld.com/article/3714621/how-llms-can-help-streamline-business-processes.html, conversational process modelling [13], and explainability of business process outcomes [6]. To date, neither there is work on leveraging LLMs for causal reasoning about business processes, nor a relevant benchmark for testing LLMs. As an early member of this family of solutions, the PET dataset [2] presents a benchmark for question answering, featuring an initial corpus of business process descriptions annotated with activities, gateways, actors, and flow information. The dataset contains 45 documents with narrative descriptions of business processes and their annotations. In [16] the TORQUE benchmark for temporal ordering has been investigated. The dataset encompasses 21k user-generated and fully answered temporal relation questions. Concerning both benchmarks, ours differs in tackling the task of causal reasoning about business processes.
@@ 4 Evaluation
We employed the developed benchmark prototype to get a sense of its applicability against state-of-the-art LLMs, to get an initial measure of accuracy spanning the three perspectives: process, causal, and the combination of the two.
At first, we ran the three core sets of template questions from Table 3 with five different target LLMs: GPT3.5444https://beta.openai.com/docs/models/gpt-3, GPT4555https://www.openai.com/research/gpt-4, GPT4o666https://www.openai.com/research/gpt-4, Mixtral-instruct-8x-7bfootnote 3, and Merlinite-7b777https://huggingface.co/ibm/merlinite-7b. We repeated each template question ten times with clean-slate prompts, preceding each question with the corresponding phrase describing the situation and instruction as illustrated in Table 6. We averaged the proportion of correctly answered questions for each reasoning perspective: process, causal, and combined. We then also added the extended set of questions (see section 3.1) and revised the results accordingly.
@@
4.1 Assessing the quality of the Benchmark
We acknowledge that our current evaluation caters strictly to the applicability of the benchmark to a handful of LLMs. We recognize that such an assessment lacks characterizing the benchmark quality. For this, we elaborate here on a set of relevant metrics that we intend to assess once the benchmark is developed in full scale.
@@ 5 Results
We report the results for running our benchmark prototype in Table 7. This table shows in percentage the proportion of questions answered correctly by each of the LLMs included in the text. The results are split between the template questions (i.e., domain-independent) and domain-specific questions, partitioned also by the various perspectives, considering the causal one without and with the addition of the extending questions. While the core set of template questions can be seen as exhaustive with respect to domain coverage, the domain-specific part should be observed with caution, as each template question was populated only with one corresponding domain-specific question. As such, we expect that in a full-scale set of domain-specific questions, the performance is likely to get closer to the accuracy presented by the template questions.
@@ 6 Conclusion and Future Work
The benchmark can be used in two ways: testing LLM performance, as demonstrated here, and training an LLM to improve its reasoning about BPC. For testing, question answers serve as the “ground truth” to determine model accuracy, which can be analyzed by perspective and domain. For training, the questions and answers can be randomly split into training and testing subsets (e.g., 80/20). The testing subset is used before and after training to measure improvement.
Setting aside the debate whether LLMs can generally reason, our goal here is more modest. We would like to equip (or measure) LLMs regarding their capacity to adequately infer sound conclusions when presented with knowledge about causal business processes. In this regard, our developed instrumentation can be employed in two (complementary) manners. It could be used to test such an ability with respect to a relatively wide variety of process domains. In addition, the tool can also be used as a model training dataset to augment an existing LLM with such ability and also to be adopted to additional problem domains for specific needs. For the former purpose, it could be used “as is” with only the questions component to facilitate model bench-marking, and with the ground truth answers for model training. For any newly embarked problem domains, the core seed of the template questions should be instantiated methodologically in the same process reported here to derive a corresponding set of domain-relevant items, to be added. This is also the rationale underlying our aim to release the model as open source (https://github.com/IBM/SAX/tree/main/NLP4BPM2024), letting the community gradually contribute to the dataset to make its domain coverage broader.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used for boosting organizational efficiency and automating tasks. While not originally designed for complex cognitive processes, recent efforts have further extended to employ LLMs in activities such as reasoning, planning, and decision-making. In business processes, such abilities could be invaluable for leveraging on the massive corpora LLMs have been trained on for gaining deep understanding of such processes. In this work, we plant the seeds for the development of a benchmark to assess the ability of LLMs to reason about causal and process perspectives of business operations. We refer to this view as Causally-augmented Business Processes (BPC). The core of the benchmark comprises a set of BPC related situations, a set of questions about these situations, and a set of deductive rules employed to systematically resolve the ground truth answers to these questions. Also with the power of LLMs, the seed is then instantiated into a larger-scale set of domain-specific situations and questions. Reasoning on BPC is of crucial importance for process interventions and process improvement. Our benchmark, accessible at https://huggingface.co/datasets/ibm/BPC, can be used in one of two possible modalities: testing the performance of any target LLM and training an LLM to advance its capability to reason about BPC.
@@
While reasoning may not be inherent in the architecture of LLMs, critics do argue that the ability to reason does manifest itself as an emergent property in LLMs [20]. That is, once trained on sufficiently large corpora of examples, LLMs may become capable of statistically deriving statements that are also coincidentally sound with the given set of arguments in the input. Similar emergent abilities may also include planning, decision-making, in-context learning, and answering in zero-shot settings [15]. Whether reasoning is an inherent capability or an emergent property of future LLMs, particularly for interpreting business process models, is likely to remain a heated debate. This debate emphasizes the importance of establishing benchmarks for evaluating LLM performance. Our focus is on driving such enablement by using LLMs to analyze textual descriptions of Causally-augmented Business Processes (BPCs). Such descriptions differ from conventional process descriptions by also including statements about the causal execution dependencies among the activities as first-class citizens. Teaching LLMs to reason about BPCs could enhance the analysis and improvement of such processes.
To standardize and measure the ability of LLMs to facilitate such tasks, a designated benchmark is being developed. The purpose of the developed novel benchmark is twofold: as a testing dataset that can be employed to quantify the ability of an LLM to reason about BPCs, and as a training dataset that can be used to adapt the ability of an LLM for this task. We evaluated two open-source and three commercial LLMs on a small subset of the situations and questions. Our results highlight the importance of producing an objective numeric scale partitioned by different perspectives to compare and assess various LLM performances. They also indicate that there is room for further improvement of LLMs in the task of reasoning about BPCs. The evolving benchmark and corresponding prompts are available here: https://github.com/IBM/SAX/tree/main/NLP4BPM2024.
@@ 2 Background
The potential of leveraging LLMs in the BPM field has been recently researched, for example, by analyzing which opportunities and challenges LLMs pose for the individual stages of the BPM lifecycle [19], in [4], where prompt engineering techniques are discussed as an alternative to fine-tuning a specific LLM, and in [12] where the authors introduce the notion of a Large Process Model (LPM), an envisioned neuro-symbolic software system that integrates process management knowledge in organizations with LLMs and statistical and inference methods for the automated inference of insights and actions. In addition, LLMs have been researched in a wide range of tasks related to BPM, including process mining [9, 3], automation of portions of complex tasks 222https://www.infoworld.com/article/3714621/how-llms-can-help-streamline-business-processes.html, conversational process modelling [13], and explainability of business process outcomes [6]. To date, neither there is work on leveraging LLMs for causal reasoning about business processes, nor a relevant benchmark for testing LLMs. As an early member of this family of solutions, the PET dataset [2] presents a benchmark for question answering, featuring an initial corpus of business process descriptions annotated with activities, gateways, actors, and flow information. The dataset contains 45 documents with narrative descriptions of business processes and their annotations. However, the PET descriptions do not include causal execution dependencies among the activities. In [16] the TORQUE benchmark for temporal ordering has been investigated. The dataset encompasses 21k user-generated and fully answered temporal relation questions. Concerning both benchmarks, ours differs in tackling the task of causal reasoning about business processes.
@@ 3.1 Extending the benchmark
Another form of extension can be achieved by extending the grammar underlying the process situations to express more complex structures (e.g., gateways).
@@ 4 Evaluation
An initial instantiation of the benchmark seed, the ‘benchmark prototype’ dataset, was populated for testing purposes, including one domain-specific question for each template question in situation #2. We used this prototype to assess its applicability and measure accuracy against state-of-the-art LLMs across three perspectives: process, causal, and their combination.
At first, we ran the three core sets of template questions from Table 3 with five different target LLMs: GPT3.5444https://beta.openai.com/docs/models/gpt-3, GPT4555https://www.openai.com/research/gpt-4, GPT4o666https://openai.com/index/hello-gpt-4o/, Mixtral-instruct-8x-7b, and Merlinite-7b777https://huggingface.co/ibm/merlinite-7b. We repeated each template question ten times with clean-slate prompts, preceding each question with the corresponding phrase describing the situation and instruction as illustrated in Table 6. We averaged the proportion of correctly answered questions for each reasoning perspective: process, causal, and combined. We then also added the extended set of questions (see section 3.1) and revised the results accordingly.
@@
We acknowledge that our current evaluation caters strictly to the applicability of the benchmark to a handful of LLMs. Such an assessment lacks characterizing the benchmark quality. For this, we elaborate here on a set of relevant metrics that we intend to assess once the benchmark is developed in full scale.
@@ 5 Results
We report the results for running our benchmark prototype in Table 7. This table shows the proportion of questions answered correctly by each LLM. The results are split between the template questions and domain-specific questions, and are also partitioned by the various perspectives, considering the causal perspective both without and with the addition of the extending questions. While the template questions provide exhaustive domain coverage, the domain-specific ones should be interpreted cautiously, as each template question has only one corresponding domain-specific question. As such, we expect that in a full-scale set of domain-specific questions, the performance is likely to get closer to the accuracy presented by the template questions. In addition, it is more likely the corpora employed for LLM training were domain-specific, hence incidentally implying an improved performance for domain-specific questions.
@@
The benchmark can be used for testing LLM performance, as demonstrated here, and for training an LLM to improve its reasoning about BPC. For testing, question answers serve as the “ground truth” to determine model accuracy, which can be analyzed by perspective and domain. For training, the questions and answers can be randomly split into training and testing subsets (e.g., 80/20). The testing subset is used before and after training to measure improvement.
@@ 6 Conclusion and Future Work
Setting aside the debate whether LLMs can generally reason, our goal here is more modest. We would like to equip (or measure) LLMs regarding their capacity to adequately infer sound conclusions when presented with knowledge about causal business processes. In this regard, our developed instrumentation can be employed in two (complementary) manners. It could be used to test such an ability with respect to a relatively wide variety of process domains. In addition, the tool can also be used as a model training dataset to augment an existing LLM with such ability and also to be adopted to additional problem domains for specific needs. For the former purpose, it could be used “as is” with only the questions component to facilitate model bench-marking, and with the ground truth answers for model training. For any newly embarked problem domains, the core seed of the template questions should be instantiated methodologically in the same process reported here to derive a corresponding set of domain-relevant items. This is also the rationale underlying our aim to release the model as open source, letting the community gradually contribute to the dataset to make its domain coverage broader. Our current version of the developed benchmark can be accessed here: https://huggingface.co/datasets/ibm/BPC.